### v2

# A Novel Vascular Risk Scoring Framework for Quantifying Sex-Specific Cerebral Perfusion from 3D pCASL MRI

Abstract
We present a novel framework that leverages 3D pseudo-continuous arterial spin labeling (pCASL) MRI to investigate sex- and age-dependent heterogeneity in cerebral perfusion and to establish a biologically informed vascular risk quantification metric. A custom convolutional neural network was trained on ASL-derived cerebral blood flow (CBF) maps from 186 cognitively healthy individuals (89 males and 97 females, ages 8-92 years), achieving 95% accuracy in sex classification and revealing robust sex-specific perfusion signatures. Regional analyses identified significantly elevated CBF in females across medial Brodmann areas 6 and 10, the visual area of the cortex, the polar occipital cortex, and both ventral and dorsal dysgranular insula, highlighting sex-specific neurovascular specialization in motor, cognitive, sensory, and affective domains. In addition, we observed a consistent global age-related decline in CBF across both sexes, reflecting progressive cerebrovascular aging. To integrate these findings, we propose a biologically informed Vascular Risk Score (VRS) derived from age- and sex-stratified normative CBF distributions. The VRS enables individualized assessment of cerebral perfusion integrity by quantifying deviations from expected normative patterns. This metric offers a sensitive, personalized biomarker for detecting early hypoperfusion and stratifying vascular contributions to neurodegenerative diseases, including Alzheimer’s disease, thereby advancing the goals of precision neurology.
Index Terms—  3D pCASL MRI, CBF, sex-specific perfusion patterns, age-based perfusion changes, vascular risk score, cognitively healthy
1 Introduction
Arterial Spin Labeling (ASL) is a non-invasive Magnetic Resonance Imaging (MRI) technique designed to quantitatively assess cerebral blood flow (CBF) by magnetically labeling endogenous arterial blood water protons. Unlike conventional perfusion imaging methods, ASL does not require exogenous contrast agents or ionizing radiation, thereby offering a safe and repeatable approach for evaluating cerebral perfusion. The technique involves the selective inversion of magnetization in arterial blood proximal to the imaging region [1], followed by acquisition of labeled and control images. The difference in signal intensity between these images reflects the delivery of magnetically tagged blood to brain tissue, enabling quantitative measurement of regional CBF.
The fundamental process of ASL MRI involves three key steps: (i) Magnetic labeling of blood: A brief radiofrequency pulse is applied to arterial blood proximal to the brain, magnetically tagging the water protons within the blood without the need for contrast agents. (ii) Inversion and delivery: The magnetically labeled blood flows into the cerebral microvasculature, delivering the inverted magnetization to brain tissue, thereby altering the local MR signal in proportion to regional perfusion. (iii) Image acquisition and subtraction: Two sets of images are acquired—a labeled (tagged) image with magnetically tagged blood and a control image without labeling. The voxel-wise subtraction of these paired images isolates the perfusion-weighted signal, enabling quantitative mapping of CBF [1].
ASL is primarily utilized to measure regional CBF, a critical biomarker in the diagnosis and characterization of numerous neurological disorders [2]. ASL enables the identification of ischemic or infarcted brain tissue in stroke patients, providing essential information regarding tissue viability and therapeutic planning. Furthermore, ASL is valuable in detecting and monitoring neurodegenerative conditions such as Alzheimer’s disease (AD), where alterations in cerebral perfusion often precede structural abnormalities detectable by conventional MRI. Additionally, ASL contributes to the evaluation of brain tumors by delineating perfusion heterogeneity and is increasingly applied in epilepsy to localize epileptogenic foci through perfusion abnormalities.
ASL offers several notable advantages over traditional CBF measurement techniques such as positron emission tomography (PET) and single-photon emission computed tomography (SPECT) [3]. These advantages include:
  • Non-invasive: ASL does not require the injection of radioactive tracers or contrast agents, thereby reducing patient risk and discomfort.
  • Quantitative: ASL enables absolute quantification of CBF, allowing for accurate evaluation of neurological conditions such as stroke, dementia, and brain tumors.
  • Repeatable: Due to its non-invasive nature, ASL can be safely repeated on the same individual, making it well-suited for longitudinal studies and monitoring treatment responses.
  • Radiation-free: In contrast to PET and SPECT, ASL uses only magnetic fields and radiofrequency pulses, avoiding exposure to ionizing radiation.
Among ASL variants, 3D pCASL MRI is the most widely adopted in both clinical and research contexts. This technique achieves an optimal balance between sensitivity [1], safety, and practical implementation, outperforming alternatives such as Continuous ASL (CASL), Pulsed ASL (PASL), and EPI-based ASL. 3D CBF maps generated via 3D pCASL provide detailed volumetric representations of cerebral perfusion [4], capturing spatial distribution and relative blood flow magnitude throughout the brain. These maps are indispensable for clinical diagnostics and research applications [5], offering critical insights into brain health and function.
CBF refers to the delivery rate of oxygen-rich blood to brain tissue and serves as a crucial indicator of neural activity and structural health [6]. As it reflects the brain’s metabolic demands and vascular integrity, alterations in CBF can signal the presence of neurological disorders [7], such as stroke, dementia, or traumatic brain injury. Reduced or asymmetrical flow may indicate ischemia or impaired autoregulation, while elevated perfusion can be associated with inflammation or hyperemia. Because of these associations, CBF is increasingly used in clinical settings to detect early pathological changes, guide therapeutic decisions, and monitor disease progression or recovery following intervention.
In the context of healthy aging, CBF can act as a sensitive biomarker of neurovascular function, as subtle declines in perfusion may precede cognitive decline or structural brain changes. Tracking CBF across the lifespan can help distinguish normal aging trajectories from early signs of neurodegeneration, thereby supporting timely preventive or therapeutic strategies.
Since 3D pCASL is primarily a structural and functional imaging technique for quantifying CBF [5], it has been widely employed to investigate sex-related differences in brain physiology and pathology [8]. Such differences in cerebral perfusion have been documented across various neurological and psychiatric disorders. The high spatial resolution and comprehensive brain coverage of 3D pCASL enable sensitive detection and mapping of these perfusion disparities through the following mechanisms:
  • Sex-related baseline differences in CBF: Studies have shown that there can be sex-based differences [8, 9] in the baseline CBF. On average, women tend to have slightly higher CBF than men. This difference has been attributed to a variety of factors, including hormonal differences (e.g., estrogen’s effects on vasodilation) and metabolic differences between males and females.
  • Age-related changes in CBF: Some studies have suggested that the rate of decline in CBF with aging may differ between men and women [9, 8]. Women may show a more pronounced decline in CBF with age, especially in certain brain regions, such as the hippocampus, which is vital for memory and is often affected in AD. Using 3D pCASL, it’s possible to observe these age-related changes in a more comprehensive manner, capturing perfusion changes in both the cortex and subcortical regions over time. Studies have shown a greater perfusion in healthy children and teenagers [10, 11]. As age increases, the perfusion rate decreases.
  • Sex differences in neurological diseases: 3D pCASL is a valuable tool to investigate how sex influences the pathophysiology of neurological diseases. For example, in AD, women tend to exhibit greater reductions in CBF compared to men [12], particularly in the parietal and temporal lobes, which are commonly affected in AD. This could reflect sex-specific differences in disease progression or pathophysiology [3]. In Parkinson’s disease, while the disease itself affects both sexes, men tend to have more severe motor symptoms at diagnosis. Differences in perfusion patterns detected via 3D pCASL could provide insights into how sex-related characteristics influence the progression of motor and non-motor symptoms.
Perfusion rates can be assessed by quantifying CBF maps derived from 3D pCASL MRI using computational algorithms. The quantification process employs a model-based approach to translate ASL signal differences into physiologically meaningful units. After acquiring paired labeled (tag) and control images [13], their subtraction yields a perfusion-weighted image reflecting the net inflow of magnetically labeled arterial blood into brain tissue during the post-labeling delay (PLD).
To obtain absolute CBF values, the perfusion-weighted signal difference is normalized by a separately acquired M0 reference image, which represents the equilibrium longitudinal magnetization of brain tissue. The M0 image [14], typically acquired via a proton density-weighted sequence with long repetition time (TR) and minimal T1 weighting, corrects for coil sensitivity variations, scanner gain, and other system factors, enabling accurate scaling of the ASL signal. Utilization of 3D acquisition sequences, such as 3D GRASE (Gradient and Spin Echo) or 3D fast spin echo, enhances signal-to-noise ratio (SNR) and spatial coverage, facilitating robust whole-brain CBF quantification.
Quantification based on the Buxton general kinetic model [15] incorporates critical parameters including labeling duration, PLD, labeling efficiency, blood–brain partition coefficient, and T1 relaxation time of arterial blood. This model assumes instantaneous delivery and retention of labeled spins within the imaging voxel and often applies a single-compartment framework, disregarding arterial transit time dispersion. Post-processing corrections—such as for partial volume effects, patient motion, and magnetic field inhomogeneities—are routinely applied to enhance accuracy. The final CBF map provides a 3D representation of perfusion, typically expressed in mL/100g/min, serving as a non-invasive quantitative biomarker for cerebral vascular health and pathology.
In standard workflows, CBF quantification from 3D pCASL images is typically performed using various neuroimaging toolkits such as BASIL, ASLtbx, ExploreASL, or ASL-MRICloud. These pipelines implement the Buxton model and automate preprocessing steps like motion correction, M0 calibration, and registration. Voxel-wise CBF values are often averaged over ROIs or anatomical atlases, with optional partial volume correction using structural MRI. While these methods are robust and interpretable, they may miss finer regional patterns in perfusion, which can be better captured by advanced approaches like supervoxel-based clustering and deep learning (DL).
Building on the capability of 3D pCASL to capture subtle perfusion differences, we propose a novel framework for the analysis of 3D CBF maps acquired via 3D pCASL MRI. Our method employs the 3D supervoxel clustering algorithm to segment brain volume into regions of homogeneous perfusion, characterized by mean intensity values extracted from each supervoxel cluster. This approach allows for robust regional characterization of cerebral perfusion heterogeneity.
Quantifying CBF from 3D pCASL data poses unique challenges [16] due to its inherently low signal-to-noise ratio (SNR), susceptibility to motion, and the spatial heterogeneity of perfusion across brain regions [17]. To address these issues, we incorporated a three-dimensional Simple Linear Iterative Clustering (3D SLIC) algorithm for supervoxel-based feature extraction. SLIC clusters adjacent voxels into compact, spatially coherent, and intensity-homogeneous regions, allowing localized CBF measurements to be aggregated into anatomically relevant and noise-robust supervoxels. This strategy enhances the reliability of CBF quantification by minimizing voxel-level noise and partial volume effects, while preserving fine-grained spatial patterns of perfusion.
In contrast to conventional region-of-interest (ROI) or voxel-wise analyses, SLIC provides a flexible, data-driven method that adapts to the local intensity structure of each individual’s brain. This reduces the feature dimensionality and increases interpretability, making it particularly suitable for DL applications where generalizability and biological relevance are critical. Furthermore, the supervoxel approach improves computational efficiency and reduces the risk of overfitting in small to moderate-sized cohorts.
Using these supervoxel features, investigate perfusion differences between cognitively normal male and female participants through a customized DL classifier. Given extensive research on aging and neurodegenerative disorders [1, 3, 12], an automated, DL-assisted sex classification system based on CBF maps could streamline diagnostic workflows, reduce clinician workload, and allow for more personalized assessments.
DL models, with their ability to learn complex feature representations [18], offer promising avenues for a rapid and accurate classification of brain perfusion patterns [19, 20]. Identifying sex-specific cerebral perfusion profiles is critical, as sex differences may influence cognitive function, neuroplasticity, and disease susceptibility. Furthermore, many neuroimaging studies currently overlook sex as a confounding variable in CBF analysis [21]. By leveraging 3D CBF maps for prediction of sex, this work aims to deepen understanding of neurovascular differences and their impact on brain function and therapeutic outcomes, ultimately improving the precision of neuroimaging-based diagnostics and research.
2 Related Works
ASL MRI has emerged as a valuable, non-invasive method for visualizing and quantifying CBF across a range of clinical and research applications. However, traditional ASL techniques face several limitations, including low signal-to-noise ratio (SNR), sensitivity to motion, and lack of vessel-selective information. Recent advances have aimed to overcome these limitations and improve the accuracy and interpretability of CBF measurements [22, 23].
One such advancement is diffusion-prepared pCASL (DP-pCASL), which incorporates diffusion weighting into the ASL process to suppress intravascular signals and enhance the specificity of tissue-level CBF measurements [24]. By reducing contamination from arterial transit artifacts and vascular components, DP-pCASL enables more accurate quantification of perfusion, particularly useful for detecting microvascular abnormalities in stroke, tumors, and neurodegenerative diseases.
Building on this enhanced specificity, DP-pCASL has proven especially effective in investigating age- and sex-related differences in cerebral perfusion. For example, Shao et al. (2024) used DP‑pCASL to quantify age- and sex-dependent trajectories of CBF, arterial transit time (ATT), and blood-brain barrier (BBB) water exchange in 186 cognitively normal participants—the same cohort studied here [8]. Their findings showed significantly higher CBF and shorter ATT in females across all age groups, underscoring robust sex differences in cerebrovascular physiology (e.g., CBF  60.5±10.7 vs. 51.5±11.9 ml/100g/min in young adults).
These observations are supported by developmental studies, such as Satterthwaite et al. (2014), which found that post-puberty, CBF declines in males but stabilizes or increases in females, particularly in networks linked to cognitive function like the default mode and executive control networks [25]. Hormonal influences—such as estrogen’s effect on vasodilation via eNOS and prostacyclin pathways—are known to elevate regional CBF in premenopausal women [26, 27, 28, 29], and structural differences, including higher gray-to-white matter ratios in females, further explain localized perfusion increases [30]. These age- and sex-specific trends in CBF are not only biologically significant but also clinically relevant, as they may signal differential vulnerability to disorders like AD, where perfusion abnormalities often precede structural degeneration [31, 32].
While DP-pCASL improves the physiological relevance of CBF measurements, another key development addresses a different limitation: the lack of vessel-selective information in standard ASL. To this end, vessel-encoded pCASL (VEPCASL) has been introduced as a technique capable of generating CBF and bolus arrival time maps for each of the four major brain-feeding arteries [33] by applying multiple postlabeling delays. In simulations, VEPCASL avoided up to 37% underestimation of CBF in voxels with dual arterial supply—a limitation inherent to conventional pCASL. In healthy volunteers, VEPCASL achieved comparable SNR and showed no systematic bias, while offering the added benefit of vessel-specific perfusion mapping. This makes it especially valuable for assessing collateral circulation in vascular pathologies.
Despite these advances, ASL-based CBF quantification remains challenging due to noise, motion artifacts, and regional variability in transit delays and labeling efficiency [16, 17]. Current software packages like BASIL and ExploreASL provide voxel-wise perfusion estimates but may struggle to account for spatial heterogeneity and can overestimate CBF in vascular-rich regions [34, 35].
To address these limitations and facilitate the integration of DL with clinically relevant perfusion metrics, we implemented a 3D SLIC) supervoxel algorithm [22] to segment CBF maps into spatially contiguous, intensity-homogeneous regions. This method enables the extraction of biologically meaningful regional features, enhances signal robustness by mitigating voxel-level noise, and reduces the dimensionality of input data for subsequent machine learning (ML) analysis. While supervoxel-based techniques have been previously employed in various MRI applications—such as prostate tissue characterization and brain tumor segmentation [36, 37, 38, 39]—to the best of our knowledge, this is the first application of SLIC-based preprocessing within the context of ASL-derived perfusion imaging.
Finally, by integrating these interpretable supervoxel-derived features into a DL framework, our method enables effective sex classification and vascular profiling. This not only provides insights into sex-specific cerebrovascular physiology but also lays the groundwork for precision neuroimaging biomarkers applicable to both clinical diagnosis and therapeutic monitoring.
3 Methodology
3.1 Participants
In this study, a publicly available dataset from OpenNeuro comprising 186 cognitively healthy participants was utilized. The cohort includes 89 males and 97 females, with ages ranging from 8 to 92 years. They are of white/Caucasian, Latinx, African American, and Asian origin.
Each participant underwent DP-pCASL MRI, from which 3D CBF maps were generated. These maps were analyzed to investigate sex classification, under the hypothesis that regional variations in mean perfusion intensity may follow sex-specific trajectories across the lifespan.
Fig. 1: Distribution of participants in each age group
Figure 1 illustrates the distribution of participants across the defined age groups, providing a visual representation of the demographic composition of the study sample.
3.2 MR data acquisition
The 3D CBF images acquired using DP-pCASL MRI were obtained from the OpenNeuro repository. These data were originally collected by Shao et al. (2024) as part of their investigation into the functional integrity of the blood-brain barrier and its variation across age and sex. Imaging was conducted on 3T Siemens Prisma scanners using either 32- or 64-channel head coils [8]. DP-pCASL acquisition parameters included a spatial resolution of 3.5 × 3.5 × 8 mm3, repetition time (TR) of 4.2 s, echo time (TE) of 36.2 ms, and a field of view (FOV) of 224 mm, with 12 slices and an additional 10% oversampling. ASL was performed with a labeling duration of 1.5 s. For a post-labeling delay (PLD) of 0.9 s, measurements were acquired at b = 0 and 14 s/mm2 across 15 repetitions. For a PLD of 1.8 s, measurements were obtained at b = 0 and 50 s/mm2 over 20 repetitions. The total scan time was approximately 10 minutes. T1-weighted structural images were acquired using 3D magnetization-prepared rapid gradient echo (MPRAGE) sequences to enable segmentation and coregistration. MPRAGE parameters included a TR of 1.6 s, inversion time (TI) of 0.95 s, TE of 3 ms, and isotropic spatial resolution of 1 mm3, with a total acquisition time of approximately 6 minutes. Slight variations in these parameters occurred across participating sites.
3.3 Proposed technique
LoadASL-derived3D cerebralblood flowmapsApplysupervoxelclusteringalgorithmExtractmeanintensitiesfromsupervoxelclustersClassify sexbased onmeanintensitiesusingdeep learningIdentify ROIsshowingperfusionpatternsdifferentlyin malesand femalesDefinesex- andage-specificvascularrisk score
Fig. 2: Proposed framework to obtain sex-based perfusion differences using 3D CBF maps
Figure 2 and Figure 3 depict the architecture of the proposed pipeline for classification of participants into males and females using cerebral perfusion imaging. The workflow consists of multiple stages designed to extract and model biologically meaningful features from 3D DP-pCASL CBF maps.
The pipeline begins with data loading and preprocessing, where raw 3D CBF maps are spatially normalized and intensity standardized to ensure inter-participant comparability. Following this, a supervoxel-based clustering technique is applied to segment the brain volume into spatially coherent regions. These supervoxels serve as data-driven regions of interest (ROIs) that preserve anatomical and functional locality while reducing dimensionality.
Within each supervoxel, mean CBF intensity is computed, resulting in a feature vector that characterizes regional cerebral perfusion for each volunteer. These features are then used to train a DL model—specifically, a neural network tailored for classification—designed to distinguish between male and female participants based on perfusion profiles.
After training, the model is evaluated on held-out data to assess classification performance using standard metrics. Finally, the pipeline includes an analysis module that investigates group-level differences in regional perfusion. This includes comparisons between male and female volunteers as well as age-stratified subgroups, allowing for the exploration of how perfusion patterns evolve across the lifespan and vary by sex. Then we analyze the variation in perfusion patterns between men and women as they age, then use the information obtained to develop a vascular risk score (VRS) that determines the vascular risk.
Fig. 3: Different stages of the proposed workflow: Stage 1—Generation of 3D CBF maps; Stage 2-Registration of CBF maps to MNI space; Stage 3—Quantification of CBF maps using 3D SLIC; Stage 4—Classification of sex, identification of significant ROIs, statistical validation and definition of VRS denoting cerebrovascular risk
This comprehensive approach not only aims to achieve accurate classification of sex but also enhances our understanding of the neurobiological differences reflected in CBF patterns.
The key contributions of this study are as follows:
  • Development of a novel CBF analysis approach - the supervoxel clustering method - that enables simplified and robust analysis of perfusion data;
  • Design and implementation of a deep learning-based sex classification framework that demonstrates high predictive accuracy using 3D DP-pCASL CBF maps;
  • Empirical characterization of sex-related differences in cerebral perfusion patterns across the brain;
  • Investigation of age-related variations in regional CBF, highlighting the interaction between aging and perfusion dynamics;
  • Designing a concept called, “Vascular risk score” to identify the risk of acquiring vascular diseases in males and females.
3.3.1 Loading of 3D CBF maps
The analysis begins with the loading of CBF data stored in NIfTI (Neuroimaging Informatics Technology Initiative) format, where each file represents a volumetric brain scan. These 3D CBF maps undergo preprocessing, including intensity normalization, to standardize the data across participants and ensure compatibility with the neural network input requirements. This step is essential for reducing inter-participant variability and enabling the model to learn consistent patterns associated with sex-specific perfusion characteristics.
Fig. 4: Colored 3D CBF maps where the upper row represents five female participants and the lower row represents five male participants; the bright red colored regions indicate increased perfusion whereas, the dark blue colored regions indicate reduced perfusion, demonstrating variability in regional cerebral perfusion patterns in both sexes
The 3D CBF maps shown in Figure 4 generated from five female and five male participants serve as visual representations of the sex-related variability in regional brain perfusion. The maps showing highly varying perfusion characteristics between male and female participants from the cohort were chosen for representation. These maps reveal pronounced spatial heterogeneities, with specific brain regions exhibiting differential perfusion intensities between males and females. In female participants, certain cortical and subcortical areas display consistently higher CBF values, which may reflect underlying neurovascular or hormonal influences. In contrast, male participants tend to show lower perfusion in corresponding regions, highlighting a pattern that persists across multiple individuals. These inter-individual and inter-sex differences in perfusion patterns are not uniformly distributed but are instead localized to particular regions.
3.3.2 Quantification of CBF
To ensure anatomical consistency across participants and enable robust group-level analyses, all CBF maps were spatially normalized to the MNI152 standard brain template (Montreal Neurological Institute) using an indirect registration approach via T1-weighted anatomical images. This process was implemented using the FMRIB Software Library (FSL). Linear (affine) registration was employed to align each subject’s CBF map—via the intermediate T2-weighted image—into a common stereotactic space. This alignment enables voxel-wise comparisons of perfusion values across individuals and supports the identification of spatially consistent CBF patterns within anatomically homologous regions [14].
Following spatial normalization, quantitative perfusion analysis was conducted through a standardized processing pipeline, incorporating key steps for accurate estimation and summarization of regional CBF characteristics. These steps are detailed in the subsequent section.
Supervoxel clustering
Supervoxel clustering is employed to segment each 3D CBF map into spatially contiguous and intensity-homogeneous regions, referred to as supervoxels. This technique extends the well-established concept of superpixels from two-dimensional image processing to volumetric neuroimaging data, enabling structured analysis of complex anatomical patterns. In this study, the 3D Simple Linear Iterative Clustering (SLIC) algorithm is utilized, which applies a modified k-means clustering approach in a combined feature space defined by voxel spatial coordinates (x, y, z) and intensity values [38]. Through iterative refinement, SLIC produces compact and approximately uniform clusters that respect intrinsic anatomical boundaries while preserving perfusion contrast [40].
Supervoxel clustering has emerged as a powerful technique in medical imaging analysis [36, 37, 38, 39, 41], offering a structured approach to capture spatial and functional heterogeneity within volumetric data. Its application enhances the accuracy and reliability of perfusion assessment by leveraging voxel similarity and spatial coherence. The key advantages of supervoxel clustering in medical data analysis include:
  • Enhanced Spatial Coherence: Supervoxel clustering aggregates voxels exhibiting similar signal characteristics, thereby preserving spatial contiguity and enabling the identification of biologically relevant regional perfusion patterns. This grouping minimizes the influence of random spatial noise and enhances interpretability of functional heterogeneity within tissue.
  • Noise Suppression: By averaging voxel-level measurements within each supervoxel, the technique effectively improves the signal-to-noise ratio (SNR). This noise reduction facilitates the detection of subtle perfusion changes that may be critical for early diagnosis and monitoring of pathological conditions.
  • Improved Regional Segmentation: Supervoxel clustering aids in the precise delineation of functional subregions characterized by distinct perfusion properties, such as differentiating ischemic tissue from normally perfused regions. This refined segmentation enhances the accuracy of quantitative assessments and supports targeted therapeutic interventions.
Applied to CBF maps generated from 3D pCASL MRI, supervoxel clustering serves to reduce data complexity and suppress voxel-level noise—both common limitations in ASL data—by aggregating voxels with similar perfusion properties into coherent regions. The brain volume of each participant is segmented into 100 supervoxels and the mean CBF intensity is calculated within each supervoxel, resulting in a regionalized perfusion profile. This process is applied uniformly across all participants, generating standardized feature representations for both male and female participants.
These supervoxels are treated as data-driven regions of interest (ROIs), enabling a biologically meaningful and computationally efficient quantification of regional cerebral perfusion. By transforming high-dimensional voxel-wise data into a compact set of interpretable features, this approach facilitates robust inter-participant comparisons and statistical analyses. Importantly, the extracted supervoxel-wise mean intensities serve as input features for downstream ML models, including sex classification tasks. Furthermore, this framework allows for systematic investigation of sex- and age-related variations in perfusion by supporting region-wise group comparisons and exploratory mapping of perfusion heterogeneity across the brain.
Extraction of mean intensities of supervoxel clusters
The primary objective of this step is to extract spatially localized characteristics of cerebral perfusion that allow meaningful comparisons between individual participants and between the male and female groups. To achieve this, the 3D SLIC algorithm is employed to partition each normalized 3D CBF map into a fixed number of supervoxels—compact, spatially contiguous regions characterized by homogeneous intensity values. This data-driven parcellation approach offers a more flexible and anatomically agnostic alternative to traditional atlas-based region definitions, capturing localized perfusion variations with higher sensitivity.
For each participant, the CBF volume is segmented into 100 supervoxels, denoted as “Cluster 1” through “Cluster 100.” Within each supervoxel, the mean intensity is computed, resulting in a 100-dimensional feature vector that characterizes the regional perfusion profile of the participant. These intensity features are subsequently associated with the participant’s sex (male or female) and grouped accordingly to facilitate sex-specific analysis of cerebral perfusion patterns.
The compactness and smoothing size were chosen to be 10 and 1 respectively, in order to maintain cross-participant consistency. The choice of 100 supervoxels was empirical, and used for the meaningful, standard, and uniform representation of participants. It helps in leveraging ASL characteristics: continuity in space and intensity. Our purpose required 100 supervoxels to identify the significant ROIs, across participants where intensities vary. More number of segments can be used according to the application.
This procedure ensures that the same number of features is extracted from each participant, enabling standardized input to downstream DL models. Importantly, all supervoxels, including those with zero and near-zero mean intensities, are retained. This inclusion guarantees consistent feature dimensionality across participants, which is critical for supervised classification tasks and statistical comparisons.
Overall, this supervoxel-based quantification framework provides a scalable, reproducible, and interpretable method for capturing localized perfusion characteristics from CBF maps, thereby supporting fine-grained analysis of physiological variability related to sex and other demographic factors.
Extraction of mean intensities of the neighboring voxels
To capture the spatial context of cerebral perfusion beyond discrete regions of interest (ROIs), we quantified the average CBF intensity within the immediate neighborhoods surrounding the supervoxel clusters identified via SLIC segmentation. While conventional neuroimaging analyses often focus exclusively on ROI-centric metrics, the perfusion characteristics of adjacent tissue provide critical insights into the vascular microenvironment, particularly in studies of cerebral aging and vascular health where interactions between microvasculature and macrovasculature may be pivotal.
For each ROI, defined by its bounding box and centroid coordinates, concentric neighborhoods were delineated by applying increasing margins at radial distances of 0.2 mm, 0.5 mm, 1 mm, and 5 mm from the centroid (assuming isotropic voxel resolution of 1 mm3). A masking procedure was implemented to isolate these peri-regional zones while excluding the core ROI itself, ensuring that measured intensities reflect the perfusion of the immediately adjacent tissue rather than the central cluster. Mean CBF intensities within these surrounding regions were computed individually for all ROIs and participants, yielding spatially resolved perfusion metrics at multiple scales.
Aggregating these neighborhood perfusion values by sex enabled direct comparisons of peri-regional blood flow patterns between male and female participants. This approach is grounded in the hypothesis that sex-specific differences in cerebral microvasculature and macrovasculature—such as capillary density, vessel diameter, and autoregulatory capacity—may manifest not only within focal perfusion hotspots but also in their surrounding vascular territories [42, 43]. By extending analysis into these perivascular zones, we aim to capture subtle sex-related variations in vascular architecture and function that influence local hemodynamics.
Our analyses revealed two notable trends:
  • A monotonic decrease in mean neighborhood perfusion intensity as a function of distance from the ROI centroid, consistent with the expected spatial gradient from macrovascular inflow regions to the surrounding microvascular beds.
  • Slightly elevated mean neighborhood intensities in females compared to males across all distances, suggesting sex-dependent differences in perfusion extending beyond primary ROIs.
These findings are consistent with prior literature demonstrating that CBF is generally higher in females compared to males across the lifespan. Numerous studies have reported that this sex-related perfusion difference is particularly pronounced in specific cortical and subcortical regions, such as the temporal and parietal lobes, which are involved in cognitive and emotional processing [8, 4]. These regional variations are believed to reflect the combined effects of hormonal modulation—particularly estrogen-mediated vasodilation—and structural differences, such as higher gray-to-white matter ratios observed in females [25, 26, 27]. Consequently, the spatial heterogeneity of perfusion observed in our results aligns with these physiological and anatomical distinctions, further supporting the use of CBF as a sensitive biomarker for studying sex-specific neurovascular characteristics.
These findings underscore the utility of neighborhood intensity metrics as potential imaging biomarkers sensitive to early vascular alterations. Changes in perivascular perfusion can precede measurable deficits within the core regions, providing a window into the pathophysiological progression of cerebrovascular and neurodegenerative disorders. In addition, the incorporation of multiple spatial scales in 3D neighborhood analyses improves the sensitivity and specificity of perfusion characterization, facilitating a more comprehensive understanding of the interaction between macrovascular supply and microvascular tissue perfusion.
3.4 Classification of participants into male and female groups using a Convolutional Neural Network (CNN)
In this study, we implemented a DL framework to classify participant sex based on regional CBF features extracted from 3D pCASL MRI data. The input features to the model consisted of supervoxel-based mean intensity values derived from CBF maps. This data-driven parcellation enables the model to capture subtle and spatially distributed perfusion patterns potentially associated with sex differences.
The perfusion disparities among men and women demonstrated in Figure 3 constitute the prominent features used by the Convolutional Neural Network (CNN) to enable robust sex classification. The CNN architecture was specifically tailored to exploit the spatial relationships inherent in the supervoxel feature vectors. Initial convolutional layers were employed to learn localized patterns of perfusion heterogeneity, preserving spatial contiguity across clusters. These layers were followed by fully connected layers designed to integrate the learned features and perform nonlinear transformations culminating in a binary classification output corresponding to biological sex.
To robustly evaluate model performance and minimize bias arising from dataset imbalances, we employed stratified 5-fold cross-validation, ensuring representative distributions of male and female participants in each fold. Throughout training, performance metrics including loss and classification accuracy were systematically tracked. This approach enabled us to assess the discriminative power of supervoxel-based CBF features in delineating sex-specific cerebral perfusion signatures. The classification results provide insights into the feasibility of employing advanced DL techniques on physiological imaging data for sex prediction, with potential implications for personalized diagnostics and understanding sex-dependent brain function.
Table 1: Summary of the CNN model
  Component Details
Input Superpixel features from CBF maps
CNN Architecture 2 Conv + Flatten + 2 FC layers
Output Linear
Loss Function BCEWithLogitsLoss()
Optimizer AdamW
Learning rate (LR) 0.0005
LR Scheduler CosineAnnealingLR
Training Strategy 5-Fold Cross-Validation
Regularization BatchNorm
Dropout 0.5
Regularization L2
Epochs 50
Batch Size 16
Table 1 represents the CNN-based DL model we have applied on the 3D CBF maps. Using the mean intensities of supervoxel clusters of 3D CBF maps from the participants, the trained CNN achieved a classification accuracy of 95%, underscoring the discriminative power of localized perfusion patterns to differentiate male and female brains. Performance metrics obtained through 5-fold stratified cross-validation further substantiate the robustness and generalizability of the model. Across all folds, CNN consistently demonstrated high accuracy, precision, recall, and F1 scores as shown in Table 2. In particular, precision and recall values were well balanced for both sexes, indicating minimal classification bias and confirming the model’s ability to generalize across heterogeneous participant data.
These results suggest that supervoxel-based CBF features effectively encapsulate subtle yet physiologically relevant sex differences in cerebral perfusion. The CNN architecture used efficiently exploits spatial hierarchies within the perfusion data, making it a robust and reliable approach for binary classification tasks based on neuroimaging biomarkers.
Table 2: Prediction using CNN model
Precision Recall F1-Score
Female 0.94 0.97 0.95
Male 0.97 0.93 0.95
Accuracy 0.95
By capturing and learning from the intricate regional patterns in the 3D CBF maps, the CNN is able to effectively differentiate between male and female brains based on their perfusion profiles. These findings not only support the feasibility of using ML approaches for perfusion imaging-based sex classification, but also underscore the broader relevance of CBF as a biomarker for exploring sex-based neurophysiological differences.
3.5 Comparing the results from other classifiers
Initially, classification of the 3D CBF maps was performed using several conventional ML algorithms, including Logistic Regression (Linear), Support Vector Machine (SVM-using Radial Basis Function kernel), Random Forest Classifier (RFC) and XGBoost Classifier (XGBC). These classifiers were selected based on their capabilities to model linear and nonlinear relationships [44], capture complex feature interactions, and efficiently process both structured and unstructured imaging data.
Table 3 presents the classification accuracies achieved by the alternative models evaluated. Ensemble-based approaches, such as the RFC and the XGBC, demonstrated improvements in robustness and predictive performance compared to simpler baseline models. However, their accuracy remained inferior to that of the convolutional neural network (CNN)-based method. This performance gap highlights the efficacy of DL architectures, which inherently capture spatially localized features and learn hierarchical representations from volumetric neuroimaging data, especially from MR imaging of the brain [44].
Table 3: Prediction consistency across classifiers
  Classifier Accuracy
Logistic Regression (Linear) 0.79
SVM (Radial Basis Function kernel) 0.87
RF Ensemble 0.82
Ensemble Boosting 0.89
These findings affirm that classification into male and female based on CBF maps derived from pCASL MRI is feasible. Given the non-invasive nature and physiological relevance of ASL imaging, utilizing CBF maps for sex classification offers promising potential for advancing personalized neuroimaging biomarkers and understanding sex-specific cerebral perfusion differences.
3.6 Age-related variation in cerebral perfusion
CBF is known to exhibit a characteristic decline with advancing age, reflecting physiological alterations in neurovascular coupling, vascular compliance, and cerebral metabolism [45]. The highest CBF values are typically observed in pediatric populations, reflecting elevated metabolic demands during brain development. As individuals transition into adolescence and adulthood, cerebral perfusion gradually decreases, with a more pronounced decline observed in older adults, which may be indicative of age-associated vascular and neurodegenerative changes.
To quantitatively characterize age-related perfusion dynamics, we computed the mean of the mean intensity values derived from supervoxel clusters, analyzed as a function of participant age and stratified by sex. This analysis facilitates the investigation of potential sex-specific differences in CBF trajectories, which may reflect the underlying vascular aging mechanisms. Figure 5 illustrates these trends, demonstrating a gradual decline in CBF with increasing age for both sexes. Across the age span, females exhibited higher average intensity values compared to males, indicating relatively preserved perfusion. These findings may partially explain the observed sex differences in vascular aging and longevity.
Fig. 5: Effect of age on perfusion in males and females
Cerebral perfusion patterns exhibit age- and sex-related variations across the lifespan. In childhood (ages 8–12), perfusion intensities are comparable between males and females, reflecting similar stages of neurodevelopment. During adolescence (13–20 years), females demonstrate slightly higher perfusion values than males. This difference becomes more pronounced in early adulthood (21–30 years), where females consistently exhibit elevated CBF relative to their male counterparts. In middle age (31–50 years), this trend persists, with females showing greater perfusion across brain regions. Notably, in the 51–70 year age range, females continue to demonstrate higher CBF levels than males. However, in older adults (71–92 years), sex-related differences diminish, and perfusion rates converge between males and females. Overall, these findings suggest a gradual decline in perfusion with age for both sexes, with females generally maintaining higher CBF throughout most of the lifespan. The observed patterns are consistent with previous literature [8, 45, 32, 46] that reports sex-dependent changes in cerebral perfusion, which can arise from differences in hormonal influences, cardiovascular risk profiles [47], and cerebrovascular reserve capacity. Importantly, the supervoxel-based approach facilitates regional specificity in age-related perfusion analyzes, allowing the detection of localized vulnerability or resilience within the vascular architecture of the brain.
These findings contribute to a growing body of evidence underscoring the importance of considering both age and sex as key biological variables in cerebral perfusion studies. This nuanced understanding can inform the development of gender-specific perfusion biomarkers for the early detection of cerebrovascular pathology and guide personalized therapeutic interventions.
3.7 Statistical analysis
To systematically characterize sex-related differences in cerebral perfusion, we performed a rigorous statistical analysis on the supervoxel-derived mean intensity features extracted from the 3D CBF maps. Specifically, a one-way Analysis of Variance (ANOVA) was conducted independently for each supervoxel cluster to evaluate whether the mean CBF intensities significantly differ between male and female cohorts. This univariate approach enables the identification of spatially localized perfusion patterns that may contribute to sex-specific cerebral hemodynamics.
The ANOVA test assesses the null hypothesis that the mean perfusion values in each supervoxel do not differ between sexes, against the alternative hypothesis of a statistically significant difference. Prior to analysis, assumptions of homogeneity of variance and approximate normality of residuals were evaluated to validate the applicability of ANOVA. Where necessary, data transformations or nonparametric alternatives could be considered, though this was not required in the present study due to adherence to statistical assumptions.
For each supervoxel feature, a one-way ANOVA was conducted to evaluate sex-related differences in perfusion, yielding an F-statistic and an associated p-value per region. These p-values represent the probability of observing a difference at least as extreme as the one found, assuming no true group effect (null hypothesis). To prioritize regions with the most pronounced group differences, features were initially ranked by ascending uncorrected p-values. Given the large number of comparisons, a multiple testing correction method, such as the Bonferroni correction, was applied to control for the family-wise error rate. Results were interpreted based on corrected p-values, with regions considered statistically significant if the adjusted p-value was below the conventional threshold ($p<0.05$).
Notably, 17 supervoxel clusters exhibited p-values below 0.05, reflecting robust and reproducible sex-dependent variations in regional CBF. These statistically significant clusters not only corroborate previously reported sex differences in cerebral perfusion but also pinpoint specific anatomical regions warranting further neurophysiological and clinical exploration. Furthermore, these discriminative features serve as valuable inputs for subsequent ML classifiers, potentially enhancing model interpretability and predictive accuracy in classification tasks.
3.8 Identification of Regions of Interest (ROIs)
To ascribe biological and anatomical relevance to the statistically significant supervoxel clusters derived from CBF maps, we employed an integrative approach combining data-driven image segmentation with atlas-based anatomical labeling. While the SLIC algorithm partitions the volumetric CBF images into supervoxels—spatially contiguous voxel clusters exhibiting homogeneous perfusion intensities—these clusters lack intrinsic neuroanatomical annotation as they are defined solely by image intensity and spatial proximity.
To contextualize the supervoxel clusters within a standardized neuroanatomical framework, the supervoxel-segmented CBF maps were co-registered to the Brainnetome Atlas—a high-resolution, connectivity-informed parcellation mapped to the MNI152 template. This atlas offers detailed cortical and subcortical delineations, enabling anatomically meaningful interpretation of regional perfusion patterns in relation to established brain structures.
The assignment of supervoxels to anatomical ROIs was operationalized by a frequently-occurring supervoxel-labeling scheme: for each supervoxel, the constituent voxels were overlaid onto the Brainnetome atlas, and the frequency distribution of ROI labels within the cluster volume (obtained from statistical analysis) was computed. The anatomical label most frequently represented within the supervoxel was then attributed to as its corresponding ROI. This strategy ensures that each supervoxel is robustly mapped to a singular anatomical region, minimizing ambiguity from partial volume effects or spatial overlap.
This enabled quantification of perfusion metrics within anatomically defined structures. Subsequently, the dataset was stratified by biological sex, allowing for group-wise statistical comparisons using independent two-sample t-tests to identify ROIs exhibiting significant sex-dependent differences in regional perfusion.
Table 4: Top 6 anatomical ROIs that help differentiate females from males based on perfusion changes
  ROIs
Brodmann Area 6
Brodmann Area 10
Area V5/MT
Occipital Polar Cortex
Ventral Dysgranular & Granular Insula
Dorsal Dysgranular Insula
Table 4 summarizes the top six anatomical regions demonstrating statistically significant differences (p<p<0.05) in mean CBF intensity between male and female participants.
Fig. 6: Identified ROIs overlaid on the brain image
Figure 6 displays the identified regions overlaid on the brain image — Medial Area 6, Medial Area 10, Area V5 of the Visual Cortex (also known as the middle temporal area (MT)), Occipital Polar Cortex, Ventral Dysgranular and Granular Insula, and Dorsal Dysgranular Insula. These regions exhibit significant sex-related differences in regional CBF, highlighting their importance as key discriminative features in sex classification models based on perfusion data.
Medial Areas 6 and 10 are part of the Brodmann’s areas (BA), a classical cytoarchitectonic mapping of the cerebral cortex widely used to relate brain structure to function. Brodmann Area 6 includes the supplementary motor area (SMA), a region critical for motor planning, coordination, and execution, especially for complex and bilateral movements. Brodmann Area 10, located in the anterior prefrontal cortex, is among the largest and most evolutionarily advanced prefrontal regions, implicated in high-level cognitive functions such as working memory, decision-making, and goal-directed behavior. These medial frontal regions are metabolically demanding and highly sensitive to changes in cerebral perfusion [48].
Our findings show pronounced age- and sex-dependent declines in CBF intensity in these Brodmann areas, with males exhibiting a greater reduction than females [49]. This pattern aligns with established evidence of sex differences in vascular aging and neurovascular coupling, influenced by hormonal factors such as estrogen, which is known to have vasoprotective and metabolic regulatory effects. The progressive CBF decline in these areas likely reflects cumulative cerebrovascular burden that contributes to impairments in motor control and executive functions seen in aging populations. Conversely, the relatively preserved perfusion observed in females may underlie their relative resilience against early neurovascular dysfunction.
Other regions, including Area V5 (MT) and the Occipital Polar Cortex—key nodes in visual motion processing and early visual perception—also show higher perfusion in females, potentially linking to sex differences in visual attention and sensory processing. The Ventral and Dorsal Dysgranular Insula, integral to the salience network and involved in interoceptive awareness, emotion regulation, and cognitive flexibility, exhibit female-biased perfusion increases, supporting sex-specific patterns of affective and adaptive neural processing.
Together, these neurovascular differences emphasize distinct sex- and age-related vulnerabilities and adaptations across motor, cognitive, sensory, and emotional brain networks. The medial Brodmann areas 6 and 10, in particular, emerge as critical biomarkers of cerebrovascular aging and sex-specific neural health, reinforcing their significance in data-driven models for sex classification based on CBF metrics.
3.9 Defining a Vascular Risk Score based on age- and sex-specific CBF intensity
Age-related reductions in CBF have been extensively documented [50, 51] and are known to reflect vascular aging and reduced neurovascular efficiency. Moreover, sex-specific physiological differences such as variations in hormonal profiles, vascular tone, and cerebral metabolism, influence baseline CBF levels, with females generally demonstrating higher perfusion than males, especially during their reproductive period. These effects are clearly evident in our analysis (Figure 5), which shows a consistent age-related decline in mean CBF intensity, with sex differences persisting across all age groups.
These findings reinforce the importance of stratifying perfusion measurements by age and sex when evaluating cerebrovascular health. To this end, we introduce a “Vascular Risk Score (VRS)”, a biologically informed, interpretable measure of vascular health derived from age- and sex-specific normative CBF values in a healthy reference cohort.
Rather than relying on z-scoring or percentile cutoffs, we define the VRS based on the empirically observed range of typical CBF variation - specifically, the group mean ($\mu_{a}^{S}$) ± one standard deviation ($\sigma_{a}^{S}$) for each age group $a$ and sex $S\in\{M,F\}$. This choice reflects an assumption of approximate normality in the healthy population distribution, where roughly 68% of individuals are expected to fall within one standard deviation of the mean. The standard deviation captures physiological inter-individual variability and thus provides a biologically grounded range of expected perfusion values.
An individual’s VRS is determined by comparing their mean CBF to the normative range:
$\left[\mu_{a}^{S}-\sigma_{a}^{S},\ \mu_{a}^{S}+\sigma_{a}^{S}\right]$
Values within this interval are considered indicative of healthy or “normal” cerebral perfusion. In contrast, individuals whose mean CBF falls below the lower bound ($\mu_{a}^{S}-\sigma_{a}^{S}$) are flagged as being “at risk”, with the computed VRS, for cerebrovascular dysfunction. The VRS is given by how far the mean CBF intensities are from the normative.
This threshold was chosen to reflect a clinically conservative boundary: while some individuals below this range may still be asymptomatic, consistently reduced perfusion has been associated with impaired neurovascular coupling, early small vessel disease, and cognitive vulnerability.
3.9.1 Participant-specific vascular risk score
To compute the VRS for a participant of a given age and sex, we apply a biologically informed comparison between the individual’s CBF and normative values derived from a healthy reference population stratified by age group $a$ and sex $S$. Let $\text{CBF}_{\text{ind}}$ denote the individual mean CBF (e.g., measured in mL/100g/min) averaged across the brain.
The computation proceeds in the following steps:
  • Identify the participant’s age group $a$ and sex $S\in\{M,F\}$.
  • Retrieve the corresponding mean CBF ($\mu_{a}^{S}$) and standard deviation ($\sigma_{a}^{S}$) from the healthy cohort.
  • Compare $\text{CBF}_{\text{ind}}$ to the lower bound of the normative interval, defined as $\mu_{a}^{S}-\sigma_{a}^{S}$.
To compare the individual’s CBF to the lower bound, the VRS is given by the deviation of the individual’s CBF from $\mu_{a}^{S}-\sigma_{a}^{S}$. Greater this deviation, greater the vascular risk. This can also be used to define a binary classification for vascular status, based on whether the individual’s CBF falls below the normative threshold as given below:
$\text{VRS}=\begin{cases}\text{Normal,}&\text{if }\text{CBF}_{\text{ind}}\geq\mu_{a}^{S}-\sigma_{a}^{S}\\ \text{At Risk,}&\text{if }\text{CBF}_{\text{ind}}<\mu_{a}^{S}-\sigma_{a}^{S}\end{cases}$
This threshold provides a conservative criterion for identifying individuals with potentially reduced perfusion, accounting for physiological variability while minimizing false positives.
While the binary classification offers clear clinical interpretability, cerebrovascular function exists on a spectrum. To quantify the extent of deviation below the normative threshold, we define a continuous measure called, “perfusion deficit” which is given as the following:
$\text{Perfusion Deficit}=\begin{cases}{\scriptstyle\mu_{a}^{S}-\sigma_{a}^{S}-\mathrm{CBF}_{ind}},&{\scriptstyle\mathrm{CBF}_{ind}<\mu_{a}^{S}-\sigma_{a}^{S}}\\ {\scriptstyle 0},&{\scriptstyle\text{otherwise}}\end{cases}$
Here, the perfusion deficit reflects how much lower an individual’s perfusion is compared to the lower bound of the normative range. A value of zero indicates that there is no detectable deficit, while positive values quantify the magnitude of the potential vascular compromise.
To express this factor on a scale suitable for risk stratification or longitudinal monitoring, we introduce a positive scaling factor $k>0$, and define the continuous VRS as:
$\text{VRS}=k\cdot\text{Perfusion Deficit}$
The factor $k$ serves several purposes:
  • Unit normalization: Ensures the score is unitless or mapped to a consistent scale.
  • Clinical tuning: Allows alignment with empirically derived risk categories or intervention thresholds.
  • Sensitivity adjustment: Controls the responsiveness of the VRS to small deviations, which may vary by population or disease context.
In practice, $k$ may be chosen empirically based on associations between VRS and downstream outcomes, such as white matter hyperintensity burden, cognitive decline, or vascular biomarkers. This approach provides a continuous, personalized index of perfusion health that enhances sensitivity to subclinical changes and supports individualized risk assessment.
3.9.2 Visualization and interpretation
Figure 7 illustrates age- and sex-stratified mean CBF values with associated standard deviations, derived from the healthy cohort. These values define the normative range against which individual perfusion is assessed.
Fig. 7: Mean CBF intensity across age groups, stratified by sex. Females (red) and males (blue) are shown separately. Error bars represent one standard deviation from the group mean, defining the normative range used to compute the VRS
In early life (ages 8–20), both males and females exhibit high CBF values, reflecting high levels of neurovascular integrity. In the 21–40 age range, a subtle divergence begins, with males showing lower mean CBF and increased variability. This trend intensifies in the mid-to-late adulthood (41–80) stage. In the oldest group (81–92), both sexes show substantial decline, reflecting a cumulative vascular burden and a narrowing of perfusion differences between the sexes.
3.9.3 Clinical relevance and implications
The proposed VRS offers several advantages for clinical and research use:
  • Biological grounding: The use of group-level means and standard deviations ensures the score reflects real physiological variation.
  • Personalization: Adjusting for age and sex allows for individualized assessment.
  • Interpretability: The binary classification offers intuitive risk categories, while the continuous version allows finer stratification.
  • Clinical utility: The VRS can be integrated into routine ASL-MRI workflows for early identification of individuals with hypoperfusion.
  • Scalability: The score is easily computable from population data and does not require complex modeling or training datasets.
While this framework does not provide a predictive model for future disease, it serves as a practical indicator of current perfusion status and potential risk. It is particularly well suited to longitudinal monitoring or inclusion in multi-modal models that combine imaging, genetic, or cognitive markers.
This simplified yet biologically informed scoring method could help bridge the gap between population-based perfusion data and personalized vascular health assessment. Future extensions may include region-specific CBF metrics, adjustment for cardiovascular or metabolic covariates, and incorporation of longitudinal trajectories.
4 Results and Discussion
Several prior studies have investigated sex-related cerebral differences and sex classification using diverse neuroimaging modalities and computational techniques. Notably, studies have confirmed the presence of sex-specific brain characteristics identifiable through imaging biomarkers [45, 52, 12]. Extending this line of research, we leverage 3D pCASL–derived CBF maps, combined with supervoxel-based regional intensity features, to develop a robust DL pipeline for sex classification in cognitively normal individuals. In addition, we introduce a novel VRS to quantify individual cerebrovascular risk using perfusion-derived features stratified by age and sex.
The proposed framework employs a supervoxel-based CNN architecture that integrates both micro- and macrovascular perfusion patterns, achieving 95% classification accuracy. This supports the biological relevance and sensitivity of our method to underlying sex differences in cerebrovascular physiology. Among regions demonstrating significant sex-based perfusion differences, the frontal lobe (BA6, BA10), occipital lobe (Area V5, Occipital polar cortex), and insula emerged as particularly salient. The frontal lobe, involved in executive function, and the occipital lobe, which supports visual processing, exhibited consistently higher perfusion in females. The insula, a key hub for interoception and salience processing, also showed robust sex asymmetries. These findings suggest that sex differences in perfusion are functionally and anatomically grounded and may be shaped by a combination of hormonal, neurovascular, and developmental factors [45, 26, 27, 28, 30, 29].
Our results align with those of Shao et al. (2024), who analyzed the same dataset and reported higher CBF in females across age groups, particularly in the frontal and occipital lobes. These reproducible regional differences point toward stable sex-related vascular patterns and reinforce the utility of perfusion-based metrics in identifying sex-specific brain organization. Notably, such disparities have been linked to estrogen-mediated vasodilation, cerebrovascular reactivity, and gray-to-white matter ratios, which differ by sex across development and aging [25, 26].
To extract interpretable regional features, we applied the SLIC algorithm to CBF maps. While SLIC-based supervoxel methods have been widely used in various MRI applications—including tissue segmentation in T2-weighted prostate imaging [36] and brain tumor segmentation in T1-weighted MRI [37, 38]—their use in ASL data preprocessing remains unexplored. To our knowledge, this is the first study to demonstrate the feasibility of applying SLIC to ASL perfusion imaging. Our results show that SLIC effectively captures spatially coherent regions with similar perfusion profiles, enabling both high classification performance and biomarker extraction from resting-state pCASL data.
The regions identified in our ROI analysis—including medial Brodmann areas 6 and 10, Area V5, the occipital polar cortex, and the dysgranular insula—exhibit significant sex- and age-related differences in CBF. Notably, medial BA6 and BA10, involved in motor and executive functions, show greater perfusion decline in males, consistent with known sex differences [49, 48] in vascular aging and estrogen’s vasoprotective role. Sensory and affective regions such as V5 and the insula display female-biased perfusion. These patterns suggest that sex-specific CBF variation in these regions may reflect differential neurovascular aging and serve as potential biomarkers of functional decline.
We also propose a novel vascular risk score (VRS), derived from the mean CBF intensities within supervoxels and benchmarked against age- and sex-specific normative values. This score serves as a non-invasive, individualized index of vascular health. Unlike conventional vascular biomarkers [50, 51] that require structural lesion detection (e.g., white matter hyperintensities) or advanced acquisitions (e.g., cerebrovascular reactivity testing), the VRS is calculated from standard resting-state ASL scans, enhancing its clinical accessibility. By incorporating demographic stratification, the VRS can detect subtle deviations from normative perfusion patterns—differences that are often obscured by global thresholds or binary classification schemes.
Despite the promising results, several limitations warrant consideration. First, the cohort size (n = 186), although diverse in age, may limit statistical power and generalizability. Additionally, relevant biological covariates—such as hormonal status, cognitive performance, vascular risk profiles, and lifestyle factors—were not included in the analysis, which could confound perfusion estimates. Second, our analysis is restricted to pCASL-derived CBF without integrating structural or functional modalities. Although SLIC supervoxels offer spatially coherent clustering, their anatomical interpretability relies on post hoc atlas mapping, which may introduce inter-subject variability. Lastly, while 5-fold cross-validation was used to mitigate overfitting, external validation on independent datasets remains essential.
Future directions include expanding the cohort to include diverse populations, incorporating additional covariates such as hormone levels, education, and socioeconomic status [8], and integrating multimodal data (e.g., fMRI, DTI) to enhance model interpretability and classification robustness. Exploring longitudinal trajectories could help distinguish transient perfusion changes from stable sex-dependent vascular traits. Additionally, domain adaptation and transfer learning strategies may improve cross-scanner and multi-center generalizability. Beyond binary classification, developing age-stratified or multi-class models could reveal nonlinear CBF trajectories across the lifespan and provide deeper insights into developmental and degenerative processes.
From a clinical perspective, targeting perfusion alterations in regions implicated in neurodegenerative disorders such as AD may support early diagnosis and intervention. Combining ASL-derived perfusion maps with functional imaging could further elucidate sex-specific neurovascular responses to pharmacologic or behavioral therapies, ultimately contributing to more precise and personalized treatment strategies. Taken together, our DL pipeline, grounded in interpretable supervoxel-based features, offers a biologically informed framework for precision neuroimaging in both health and disease.
5 Conclusion
In this study, we present a framework leveraging 3D pCASL MRI to quantify CBF through supervoxel clustering, enabling anatomically and physiologically meaningful segmentation of brain perfusion patterns. By extracting CBF intensities from spatially contiguous supervoxels in 186 cognitively normal adults, we characterized sex- and age-specific variations in cerebral perfusion. A custom CNN trained on these supervoxel-level features demonstrated robust sex classification performance, achieving high accuracy, thereby, validating the discriminative value of regional CBF metrics as a biomarker. Using the intensities, we developed a novel participant-specific metric called, “Vascular risk score” that will reveal the risks of acquiring vascular diseases as people age.
These findings project significant regional differences in cerebral perfusion between males and females, reflecting underlying neurovascular and physiological variations relevant to healthy brain aging. The approach offers a sensitive and non-invasive biomarker framework that captures both microvascular and macrovascular contributions to brain perfusion. Importantly, this methodology has the potential to enhance understanding of sex-specific brain aging trajectories and facilitate early detection of neurodegenerative conditions such as AD.
6 Conflict of Interest Statement
All authors declare no conflicts of interest.
7 Data Availability Statement
The 3D pCASL MRI data used for the study is available in the open-source database, OpenNeuro.
8 Acknowledgements
This study was conducted as part of the Funding for Aging Brain Research in Collaboration (FABRIC) Grant, initiated by the Centre for Brain Research, Indian Institute of Science, Bengaluru (authors S.N., N.S., and T.I.), and the Indian Institute of Science, Bengaluru (author V.S.). Author V.S. is also supported by the DBT/Wellcome Trust India Alliance Fellowship [IA/E/22/1/506763] and the Council of Scientific & Industrial Research (CSIR), India, under its ASPIRE (Women Scientist Scheme) program [25WS(013)/2023-24/EMR-II/ASPIRE]. The authors thank the Director and Administration of the Centre for Brain Research, Indian Institute of Science, Bengaluru, for their support throughout the study.